OpenStudy (kj4uts):

Out of 100 students sampled, 70 of them said that they hoped to get married someday. With 99.7% confidence, what is the approximate percentage of the students in the population who hope to get married someday? A. 60.8% to 79.2% B. 65.4% to 74.6% C. 56.3% to 83.7% D. 51.7% to 88.3%

OpenStudy (kj4uts):

I have added them all up and divided by two and they all give me 70. I know that 99.7% confidence equals +- 3 I am just not sure what to do from their?

OpenStudy (kj4uts):

@jim_thompson5910

OpenStudy (kj4uts):

I know that 70 of 100 is 70% (0.70)

jimthompson5910 (jim_thompson5910):

use the formulas L = p - z*sqrt(p*q/n) U = p + z*sqrt(p*q/n) where... p = sample proportion z = critical value q = 1-p n = sample size

OpenStudy (kj4uts):

I don't know if I plugged the numbers in right .7-sqrt(.7)(1-.7)/100 and got 0.654 with I think would be 65.4%

jimthompson5910 (jim_thompson5910):

you forgot about z

OpenStudy (kj4uts):

what number is z is that 3

OpenStudy (kj4uts):

I got C. 56.3% to 83.7%

jimthompson5910 (jim_thompson5910):

correct